OpenStudy (anonymous):

Need some help please.. Use Variation of Parameters: y''-y=e^x+1 I need some help getting started with Yp please

OpenStudy (anonymous):

The problem shows as: \[y''-y=e^x+1\] I havent done these before so I would like some help for the first couple problems..

OpenStudy (anonymous):

Eh, why do we have to use variation of parameters?

OpenStudy (anonymous):

Its the technique im trying to learn..

OpenStudy (anonymous):

there is another problem: \[y''+y=cscx\]

OpenStudy (anonymous):

You have a differential equation of the form: y'' + q(x)y' + r(x)y = g(x) First, identify q(x) and r(x), and g(x).

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Would you mind helping me out for a sec? @AccessDenied

OpenStudy (anonymous):

If you're supposed to use variation of parameters, you have to find two linearly independent solutions. You'll get those from the characteristic solution.

OpenStudy (anonymous):

@jon91d, you do know how to find the general characteristic solution, don't you?

OpenStudy (anonymous):

You mean the Yc? \[Yc=C1e^{-x}+C2e^x\]

OpenStudy (anonymous):

Yes. The set up for VoP requires taking two known solutions, call them \(y_1\) and \(y_2\). In this case, \(y_1=e^{-x}\) and \(y_2=e^x\).

OpenStudy (anonymous):

Ok, gotcha, is this where you need to apply the Wronskian?

OpenStudy (anonymous):

Almost there. The solution to the equation will be of the form, \[y=u_1y_y+u_2y_2\] The formulas for the \(u\)'s involves taking the Wronskian. \[u_1=-\int\frac{y_2f(x)}{W(y_1,y_2)}~dx\\ u_2=-\int\frac{y_1f(x)}{W(y_1,y_2)}~dx\] where \(f(x)\) comes from the nonhomogeneous part of the original equation. So that would be \(f(x)=e^x+1\). The Wronskian would be the determinant, \[W(y_1,y_2)=\begin{vmatrix}y_1&y_2\\\frac{dy_1}{dx}&\frac{dy_2}{dx}\end{vmatrix}\]

OpenStudy (anonymous):

So the Wronskian would be? d